OpenStudy (anonymous):

can someone help me with #6 and #7?

OpenStudy (alexwee123):

well for the first one we can see that y=3 and also that the width is 4 so 4=3-x x=-1

OpenStudy (phi):

can you write down the equations for question 7?

OpenStudy (anonymous):

what equation?

OpenStudy (phi):

I asked first. use the idea that opposite sides of a parallelogram are equal

OpenStudy (anonymous):

ya..i know you asked first. clearly i dont know what equation you are referring to.

OpenStudy (phi):

for question 7 use the idea that opposite sides of a parallelogram are equal to write down 2 equations.

OpenStudy (anonymous):

BA =CD?

OpenStudy (phi):

yes, but use the "values" they are given

OpenStudy (anonymous):

2x=7?

OpenStudy (phi):

pretty close, but AB is not just 2x. you missed -y

OpenStudy (anonymous):

2x-y=7 ...

OpenStudy (eyust707):

correct!

OpenStudy (eyust707):

Can you write down one other equation that relates BC and AD?

OpenStudy (phi):

you need 2 equations to solve for x and y (2 unknowns x and y mean you need 2 equations) can you write down the other equation?

OpenStudy (anonymous):

x+y=2

OpenStudy (phi):

ok now we have x+y=2 2x-y= 7 it is time to do algebra do you know how to add these 2 equations together?

OpenStudy (anonymous):

2x+x?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

add both sides like this x+ y + 2x -y = 2+7 can you combine "like terms" ? if you have 1 x and add 2 x's how many x's do you get? if you have 1 y and take away 1 y, how many y's do you have?

OpenStudy (phi):

can you simplify the right hand side? 2+7 = ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

ok...lets pick up from where phi left off! so what is 1x + 2x?

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

right! and y-y?!

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

perfect! so now you have...3x = 9! so..can u find out what "x" can possibly be?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

correct!! now put that value of x in any of the equations and solve for y....so... if u want to choose the first one... u say...3 + y = 2 ...can u find out the value of y?

OpenStudy (anonymous):

3+3=2?

OpenStudy (anonymous):

come on! 3 + 3 is 6! not 2!!

OpenStudy (anonymous):

i'll give u a hint! it has to be a negative number!!

OpenStudy (anonymous):

clearly! I know that i was plugging in 3 for Y!

OpenStudy (anonymous):

u know this! lets say, you have 3 apples...and ate some, and u ended up with only two apples! so, how many did you eat?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

perfect!!! in this scenario, u ate one! mathematically, u can call this as -1! :D and there's your values for x and y!!

OpenStudy (anonymous):

thank youu! :)

OpenStudy (anonymous):

glad to help! and thanks to @phi also!

OpenStudy (anonymous):

yes!